OpenStudy (anonymous):

solve: pie r^(2)h for h

OpenStudy (anonymous):

I need to double check: where is the h? On the 'ground' with r, or in the exponent with 2?

OpenStudy (anonymous):

on the ground, r is being squared

OpenStudy (anonymous):

Awesome. so to make it look nicer, we'll write it as: pi*h*r^2

OpenStudy (anonymous):

okay that looks better

OpenStudy (anonymous):

Actually, come to think of it, we need to know something... did they say the equation = 0? Because we can't solve it without an equal sign.

OpenStudy (anonymous):

no its v=pi*h*r^2 for h?

OpenStudy (anonymous):

Ah, ok. We've gotta have that v to solve this. :P

OpenStudy (anonymous):

What we want to do is try to get the h by itself. So how do you get pi to go away?

OpenStudy (anonymous):

I must go in a couple of minutes, so lemme know soon!

OpenStudy (anonymous):

divide both sides?

OpenStudy (anonymous):

by pi?

OpenStudy (anonymous):

Yup! So what do we get?

OpenStudy (anonymous):

v/pi= r^2*h

OpenStudy (anonymous):

perfect! Now what else is 'bothering' our h?

OpenStudy (anonymous):

the r^2

OpenStudy (anonymous):

And how do we make the r^2 go to the other side?

OpenStudy (anonymous):

square root?or divide?

OpenStudy (anonymous):

what kinda math is this?

OpenStudy (anonymous):

divide, same as the pi

OpenStudy (anonymous):

@arboring: Looks like physics and algebra.

OpenStudy (anonymous):

so the answer is v/pi*r^2=h?

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No prob.